# Unveiling Complex Territorial Socio-Economic Dynamics: A Statistical Mechanics Approach

### v1

@@ Abstract
This study proposes a novel approach based on the Ising model for analyzing the observed territorial configuration of a network of municipalities classified as being central hubs or peripheral areas. This is interpreted as being a reference of a system of interacting territorial binary units. The socio-economic structure of the municipalities is synthesized into interpretable composite indices, which are further aggregated by means of Principal Components Analysis in order to reduce dimensionality and construct a univariate external field compatible with the Ising framework. Monte Carlo simulations via parallel computing are conducted adopting a Simulated Annealing variant of the classic Metropolis-Hastings algorithm. This ensures an efficient local exploration of the configuration space in the neighbourhood of to the reference of the system. Model consistency is assessed both in terms of energy stability and the likelihood of these configurations. The comparison between observed configuration and simulated ones is crucial in the analysis of multivariate phenomena, concomitantly accounting for territorial interactions. Model uncertainty in estimating the probability of each municipality being a central hub or peripheral area is quantified by adopting the model-agnostic Conformal Prediction framework which yields adaptive intervals with guaranteed coverage. The innovative use of geographical maps of the prediction intervals renders this approach an effective tool. It combines statistical mechanics, multivariate analysis and uncertainty quantification, providing a robust and interpretable framework for modeling socio-economic territorial dynamics, with potential applications in Official Statistics.
@@ 1 Introduction
The increasing availability of geo-referenced demographic and economic data constitutes a challenging opportunity in spatial data analysis. Descriptive approaches may prove to be inadequate when managing data complexity if territorial structures and socio-economic aspects are combined. Traditional spatial-econometric approaches are usually based on autocorrelation and regional dependencies, providing important state-of-the-art methodology in applied statistics [1, 2]. These models require a non-trivial construction of a pre-defined spatial weight matrix and rely on assumptions of linearity and normality of input data, which may not hold in real-world applications. As a result, they may fail to capture complex non-linear relationships as well as the spatial heterogeneity of socio-economic data, resulting in a limited generalization capability. Composite indices methodology offers a more flexible approach for analyzing the influence of diverse factors related to local units such as regions or municipalities, providing a valid alternative to the aforementioned descriptive and traditional methods. These widely used methods are adopted in official statistics to synthesize multiple dimensions of complex phenomena into a single value, enabling territorial comparisons across a set of statistical units for the evaluation of multidimensional well-being, deprivation, and social vulnerability. A most relevant example is the research work carried out by the Italian BES (equitable and sustainable well-being, Italian acronym) Committee for measuring multidimensional economic, social, and policy domains [3]. These measures facilitate the comprehension of the underlying phenomena when the rank of statistical units is analyzed with the aim of enhancing effective communication between researchers and policy makers. Various composite methods based on simple additive aggregation functions are widely adopted by the scientific community, even though the theoretical assumptions required for their application may limit their effectiveness in practical cases where a complete substitutability assumption does not hold (i.e., it is not true that a decrease in one dimension can be compensated by an increase in another). As a consequence, a non-compensatory composite index for spatial as well as spatial-temporal comparisons has been successfully proposed in the literature [4]. It is important to point out that, despite the widespread diffusion of these methods in social research, composite indices fail in describing the effect of interactions between the statistical units under consideration, proving to be inadequate for analyzing complex territorial dynamics. In the field of machine learning, spatial methods are gaining interest in official statistics, where predictive modeling of heterogeneous socio-economic data is a standard task, as in the case of convolutional neural networks for estimating poverty by analyzing satellite imagery combined with socio-economic predictors [5]. This integration of spatial statistics with data-driven approaches is particularly effective in gaining a deeper insight into social inequalities across a given territory. In the same context of territorial analysis using machine learning, a novel approach based on region-specific boosted classification trees has been proposed to evaluate the importance of a set of composite dimensions in relation to an observed classification of Italian municipalities as either central hubs or peripheral areas [6]. The results of this approach illustrate the potential of using composite indices in machine learning while considering the territory as divided into a set of disjoint partitions, without accounting for interactions between local units. These limitations suggest that models integrating local spatial interactions with external influences to interpret observed territorial arrangements of local units may prove to be more robust when compared to other classification approaches. In this perspective, this study proposes the Ising model from Statistical Mechanics, a versatile algorithm for describing interactions among particles in complex physical systems [7]. This method has also been successfully applied to opinion dynamics, biological networks, and socio-economic territorial domains [8, 9]. A key feature of this model consists in the effective integration of local interactions with external influences in a probabilistic modeling framework. In the specific context of spatial analysis, this method may describe the asymptotic behavior of a network of interacting local units subject to an external field of observed influences, by taking the adjacency of the units into account as is the case of municipalities with territorial attributes subject to socio-economic influences. In such a case the adjacency of local units is interpreted in terms of both geographical proximity and territorial similarity. Due to the combinatorial complexity of such a network, the exploration of all possible configurations of the system is not feasible. Therefore, the system is simulated by means of Markov Chain Monte Carlo (MCMC) techniques based on the Metropolis-Hastings algorithm, in order to generate configurations of binary local units sampled from an unknown probability distribution as the system reaches stationarity [10, 11]. This simulation process may be computationally intensive. To reduce the execution time and computational burden, a parallel computing approach is mandatory in most real-world applications [12, 13]. The randomly generated configurations are used to estimate the probability of each local unit being in a particular state, thereby assessing the contribution of socio-economic aspects while accounting for territorial similarities in the emergence of specific patterns in the estimated classification of municipalities. The comparison with the observed classification of municipalities is crucial for evaluating the socio-economic influences, which are defined by a set of composite indices summarizing social, demographic, and economic characteristics. Principal Components Analysis is then applied to further aggregate these indices into a single latent structure, referred to as the external field in the Ising model. The model is simulated using Simulated Annealing (SA), a variant of the aforementioned MCMC approach, aimed at iteratively exploring the configuration space by flipping nodes according to local energy changes [14, 15]. Simulations are initialized from the observed territorial classification, which serves as the reference configuration. As reported in the literature, Simulated Annealing is asymptotically convergent to a global optimum when a logarithmic cooling schedule is adopted, regardless of the initial configuration and initial temperature of the system. This study proposes an approach that diverges from this theoretical setting by employing a localized stochastic search for more energetically favorable and more likely configurations in the neighbourhood of the reference. As a consequence, an adaptation of the standard Ising framework is proposed, where the updating mechanism is modified in a stochastic gradient descent perspective. The reliability of probability estimates is assessed through the Conformal Prediction (CP) framework [16, 17, 18, 19]. The methods in this framework evaluate distribution-free prediction intervals with guaranteed coverage of the true value of the target variable of a machine learning model, even in the absence of traditional model assumptions [20, 21]. The proposed approach aims to contribute to the field of spatial analysis by combining methods from Statistical Physics, Multivariate Analysis, and Conformal Prediction. The Ising model framework is innovatively extended for socio-economic analysis. The incorporation of Conformal Prediction introduces a crucial step for quantifying uncertainty through the construction of adaptive prediction intervals, which are projected onto geographical uncertainty maps. These maps represent a valuable tool for detecting areas of limited model reliability in policy-driven territorial analysis, as highlighted by the results from the real-world case study presented in this article.

### v2

@@ Abstract
This study proposes a novel approach based on the Ising model for analyzing socio-economic emerging patterns between municipalities by investigating the observed configuration of a network of selected territorial units which are classified as being central hubs or peripheral areas. This is interpreted as being a reference of a system of interacting territorial binary units. The socio-economic structure of the municipalities is synthesized into interpretable composite indices, which are further aggregated by means of Principal Components Analysis in order to reduce dimensionality and construct a univariate external field compatible with the Ising framework. Monte Carlo simulations via parallel computing are conducted adopting a Simulated Annealing variant of the classic Metropolis-Hastings algorithm. This ensures an efficient local exploration of the configuration space in the neighbourhood of the reference of the system. Model consistency is assessed both in terms of energy stability and the likelihood of these configurations. The comparison between observed configuration and simulated ones is crucial in the analysis of multivariate phenomena, concomitantly accounting for territorial interactions. Model uncertainty in estimating the probability of each municipality being a central hub or peripheral area is quantified by adopting the model-agnostic Conformal Prediction framework which yields adaptive intervals with guaranteed coverage. The innovative use of geographical maps of the prediction intervals renders this approach an effective tool. It combines statistical mechanics, multivariate analysis and uncertainty quantification, providing a robust and interpretable framework for modeling socio-economic territorial dynamics, with potential applications in Official Statistics.
@@ 1 Introduction
The increasing availability of geo-referenced demographic and economic data constitutes a challenging opportunity in spatial data analysis. Descriptive approaches may prove to be inadequate when managing data complexity if territorial structures and socio-economic aspects are combined. Traditional spatial-econometric approaches are usually based on autocorrelation and regional dependencies, providing important state-of-the-art methodology in applied statistics [1, 2]. These models require a non-trivial construction of a pre-defined spatial weight matrix and rely on assumptions of linearity and normality of input data, which may not hold in real-world applications. As a result, they may fail to capture complex non-linear relationships as well as the spatial heterogeneity of socio-economic data, resulting in a limited generalization capability. Composite indices methodology offers a more flexible approach for analyzing the influence of diverse factors related to local units such as regions or municipalities, providing a valid alternative to the aforementioned descriptive and traditional methods. These widely used methods are adopted in official statistics to synthesize multiple dimensions of complex phenomena into a single value, enabling territorial comparisons across a set of statistical units for the evaluation of multidimensional well-being, deprivation, and social vulnerability. A most relevant example is the research work carried out by the Italian BES (equitable and sustainable well-being, Italian acronym) Committee for measuring multidimensional economic, social, and policy domains [3]. These measures support the understanding of complex socio-economic phenomena and facilitate effective communication between researchers and policy makers, especially when rankings of statistical units are analyzed. Composite indicators play a central role in the field of social indicators, offering essential tools for monitoring inequality, well-being, and regional cohesion. They are widely adopted by national statistical institutes and international frameworks to evaluate social progress and guide public policy across space [5, 6]. These indicators provide a multidimensional representation of well-being, going beyond income-based measures and enabling spatially informed decision-making. Despite their widespread use, most composite indices rely on simple additive aggregation functions. Although broadly accepted, these methods often assume complete compensability across dimensions—meaning that a deficit in one dimension can be offset by a surplus in another—which may not hold in real-world settings. To address this limitation, a non-compensatory composite index has been proposed for both spatial and spatial-temporal comparisons [4]. A critical limitation remains: composite indices, whether compensatory or not, do not capture interactions between statistical units. As such, they may prove inadequate for analyzing complex territorial dynamics that emerge from interdependencies among local units. In the field of machine learning, spatial methods are gaining interest in official statistics, where predictive modeling of heterogeneous socio-economic data is a standard task, as in the case of convolutional neural networks for estimating poverty by analyzing satellite imagery combined with socio-economic predictors [7]. This integration of spatial statistics with data-driven approaches is particularly effective in gaining a deeper insight into social inequalities across a given territory. In the same context of territorial analysis using machine learning, a novel approach based on region-specific boosted classification trees has been proposed to evaluate the importance of a set of composite dimensions in relation to an observed classification of Italian municipalities as either central hubs or peripheral areas [8]. The results of this approach illustrate the potential of using composite indices in machine learning while considering the territory as divided into a set of disjoint partitions, without accounting for interactions between local units. These limitations suggest that models integrating local spatial interactions with external influences to interpret observed territorial arrangements of local units may prove to be more robust when compared to other classification approaches. Graph-based models have recently gained increasing attention for their ability to represent spatial systems through topological rather than strictly geographical adjacency. In these approaches, spatial units are interpreted as nodes in a network, with edges encoding various forms of structural or functional similarity rather than physical contiguity. This allows for the modeling of spatial interactions in highly heterogeneous systems where classical distance-based metrics may be inadequate. In human geography and urban planning, graph representations have been employed to study socio-spatial organization, infrastructure networks, and regional dependencies [9, 10]. These models emphasize the conceptual nature of spatial relationships, offering a flexible framework for representing territorial complexity. The approach proposed in this study builds on this graph-based perspective by modeling socio-economic interactions between municipalities based on shared territorial profiles, independently of strict geographical proximity. In this perspective, this study proposes the Ising model from Statistical Mechanics, a versatile algorithm for describing interactions among particles in complex physical systems [11]. Recent studies have emphasized the potential of adopting tools from statistical physics and network science for modeling socio-economic and territorial systems. Approaches based on the Ising model and its variants have been successfully applied to understand regional disparities [12], structural dependencies in socio-economic resilience [13], and spatial complexity in urban systems [14]. These works suggest that integrating spatial interactions within probabilistic frameworks can enhance the interpretability and robustness of territorial models. This method has also been successfully applied to opinion dynamics, biological networks, and socio-economic territorial domains [15, 16]. A key feature of this model consists in the effective integration of local interactions with external influences in a probabilistic modeling framework. In the specific context of spatial analysis, this method may describe the asymptotic behavior of a network of interacting local units subject to an external field of observed influences, by taking the adjacency of the units into account as is the case of municipalities with territorial attributes subject to socio-economic influences. In such a case the adjacency of local units is interpreted in terms of both geographical proximity and territorial similarity. Due to the combinatorial complexity of such a network, the exploration of all possible configurations of the system is not feasible. Therefore, the system is simulated by means of Markov Chain Monte Carlo (MCMC) techniques based on the Metropolis-Hastings algorithm, in order to generate configurations of binary local units sampled from an unknown probability distribution as the system reaches stationarity [17, 18]. This simulation process may be computationally intensive. To reduce the execution time and computational burden, a parallel computing approach is mandatory in most real-world applications [19, 20]. The randomly generated configurations are used to estimate the probability of each local unit being in a particular state, thereby assessing the contribution of socio-economic aspects while accounting for territorial similarities in the emergence of specific patterns in the estimated classification of municipalities. The comparison with the observed classification of municipalities is crucial for evaluating the socio-economic influences, which are defined by a set of composite indices summarizing social, demographic, and economic characteristics. Principal Components Analysis is then applied to further aggregate these indices into a single latent structure, referred to as the external field in the Ising model. The model is simulated using Simulated Annealing (SA), a variant of the aforementioned MCMC approach, aimed at iteratively exploring the configuration space by flipping nodes according to local energy changes [21, 22]. Simulations are initialized from the observed territorial classification, which serves as the reference configuration. As reported in the literature, this algorithm is asymptotically convergent to a global optimum (if it exists) when a logarithmic cooling schedule is adopted, regardless of the initial configuration and initial temperature of the system. This study proposes an approach that diverges from this theoretical setting by employing a localized stochastic search for more energetically favorable and more likely configurations in the neighbourhood of the reference. As a consequence, an adaptation of the standard Ising framework is proposed, where the updating mechanism is modified in a stochastic gradient descent perspective. The reliability of probability estimates is assessed through the Conformal Prediction (CP) framework [23, 24, 25, 26]. The methods in this framework evaluate distribution-free prediction intervals with guaranteed coverage of the true value of the target variable of a machine learning model, even in the absence of traditional model assumptions [27, 28]. The proposed approach aims to contribute to the field of spatial analysis by combining methods from Statistical Physics, Multivariate Analysis, and Conformal Prediction. The Ising model framework is innovatively extended for socio-economic analysis. The incorporation of Conformal Prediction introduces a crucial step for quantifying uncertainty through the construction of adaptive prediction intervals, which are projected onto geographical uncertainty maps. These maps represent a valuable tool for detecting areas of limited model reliability in policy-driven territorial analysis, as highlighted by the results from the real-world case study presented in this article.